OpenStudy (anonymous):

What is the value of z in the equation 2(4z − 5 − 3) = 166 − 46? A) 15 B) 17 C) 30 D) 32

OpenStudy (anonymous):

@ssssss

OpenStudy (anonymous):

2(4z-8)=120

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

8z-16=120. 8z=120+16 8z=136 z=17

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thankkk youuu sooo much

OpenStudy (anonymous):

You are welcome))